OpenStudy (anonymous):

what's the derivative of (4x+9)?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

okay so I tried using the quotient rule to find the derivative of (6x)/(4x+9) and I got the wrong answer. any help?

OpenStudy (anonymous):

quotient rule dy/dx = vdu - udv -------- if y = u/v v^2 y =6x / 4x+9 y' = (4x+9)* 6 - 6x * 4 ---------------- (4x+9)^2 = 24x + 54 - 24x ------------ (4x+9)^2 = 54 --- (4x+9)^2

OpenStudy (anonymous):

hahah alright thanks. that's what i got the first time i did it, but it didn't accept my answer because i factored out (4x+9)^2. so stupid.